OpenStudy (anonymous):

What is the equation of the line, in slope-intercept form, that passes through (3, -1) and (-1, 5)? y= -2y= -3x+ 7 2x+ 3y- 7 = 0

OpenStudy (ivettef365):

first find the slope

OpenStudy (anonymous):

idk the slope

OpenStudy (ivettef365):

do you have the formula?

OpenStudy (anonymous):

no my teacher didnt give 1

OpenStudy (ivettef365):

y2 - y1 m = --------- x2 - x1

OpenStudy (ivettef365):

x1,y1 x2,y2 you have points (3,-1) and (-1,5)

OpenStudy (anonymous):

ok

OpenStudy (texaschic101):

m = (y2 - y1) / (x2 - x1)......m is the slope....set 1 (3,-1)...set 2 (-1,5) m = (5 - (-1) / (-1 - 3) m = 6/-4 = - 3/2 the slope is - 3/2 y = mx + b <--slope intercept form we will be using slope - 3/2 and points (3,-1) -1 = -3/2(3) + b -1 = - 9/2 + b -1 + 9/2 = b -2/2 + 9/2 = b 7/2 = b y = -3/2x + 7/2 <--this is your equation in slope intercept form.